Identify the surface of the z= 4 + 4r2 by converting them into equations in the Cartesian form. Show the complete solutions.



1
Expert's answer
2022-06-03T05:05:34-0400

The surface is given in cylindrical coordinates "(r, \theta, z)," and the conversion formula:

"x=r cos \theta,"

"y=r sin \theta,"

"z=z."

Since "x^2+y^2=r^2," we can convert the equation "z^2 = 4 + 4r^2" into Cartesian form:

"z^2=4+4(x^2+y^2),"

"-4x^2-4y^2+z^2=4."

By dividing the equation by 4 we obtain the equation of the hyperboloid of two sheets:

"-x^2-y^2+\frac{z^2}{4}=1,"

"-\frac{x^2}{1^2}-\frac{y^2}{1^2}+\frac{z^2}{2^2}=1."